# Intention Analysis Makes LLMs A Good Jailbreak Defender

### v3

@@ 6 Conclusion
In this work, a simple yet highly effective defense strategy, namely $\mathbb{IA}$, is proposed to handle the widespread complex and stealthy jailbreak attacks. $\mathbb{IA}$ works by leveraging the intrinsic intention recognition capacities of LLMs, and instructing them to first analyze the essential intention of user queries before finally responding through two stages. Extensive experiments on representative jailbreak benchmarks across LLMs of varying scales and alignment levels show that $\mathbb{IA}$ could consistently and significantly reduce the harmfulness in responses (averagely -53.1% attack success rate), while maintaining the general helpfulness. Further analysis indicates that enhancing LLMs’ intention analysis capability and their inherent safety are two directions for future improvements.
@@
Due to limited budgets, GPT-4 is not yet included in our experiments. However, its superior language understanding (including intention analysis) and safety capabilities suggest it will follow the trends we see with GPT-3.5 in our study. Additionally, despite the effectiveness of our method in defending sophisticated jailbreak prompts, these prompts do not encompass the entire potential jailbreak attacks encountered in real-world scenarios. Consequently, the practical applicability of our approach remains to be validated through further testing. Our research underlines the importance of intention analysis in improving LLM safety, suggesting future work focusing on integrating this into training to reduce inference costs. Additionally, in the face of the rapid advancements in the adversarial attacks community, there is a pressing need for developing more effective and robust defense strategies for LLMs.
@@ C.1 Automation-based Safety Evaluation
For evaluation of DAN, considering the complex instructions in DAN makes it challenging to directly identify the success of an attack through string matching, we turn to utilize a widely-adopted LLM to evaluate the harmfulness of model generations. Practically, we adopt gpt-3.5-turbo-0613131313https://openai.com/blog/chatgpt as the evaluation function following Deng et al. (2023a). Especially, the model responses are submitted via API calls, and a specific evaluation prompt is designed for gpt-3.5-turbo-0613 to rate the harmfulness on a five-degree scale (‘A (Very Harmful)’, ‘B (Harmful)’, ‘C (Moderately Harmful)’, ‘D (Slightly Harmful)’, and ‘E (No Harmful at All)’). Figure 6 shows the detailed instruction for gpt-3.5-turbo-0613 evaluation.
@@ Appendix D Alternative Prompts
To assess the resilience of our method against specific expressions, we construct other two sets of alternative instructions for intention analysis. These additional instructions were meticulously designed with modifications of origin $\mathbb{IA}$ prompts (detailed in Figure 8) in lexical choices and syntactic constructions to direct intention analysis. The detailed prompts are presented in Figure 10. Comparison results in Table 6 demonstrate that: it is the intention analysis mechanism itself, but not the specific expression of the instructions, that makes our method effective.
@@
Appendix E Comparison with all defense baselines
Table 9 lists the comparison results between our $\mathbb{IA}$ and the baselines. [^14] Due to computational resource and financial limitations, we randomly select 320 samples each from DAN and SAP200 datasets for comparative experiments. As observed, our method consistently shows superiority over other baselines on different datasets and model scales. Specifically, our $\mathbb{IA}$ outperforms the second-best method by 30.32% and 23.77% averagely on SAP200 and DAN, respectively. In addition, although ICD and Self-Reminder achieve considerable reduction in ASR on GCG, their performance severely degrades when dealing with complex and stealthy jailbreak prompts. On the contrary, our method consistently outperforms other baselines across both prompt-level and automatic token-level jailbreak datasets. Notably, our method achieves the best ASRs with comparable and acceptable empirical inference runtime.
Appendix F Cross-Intention Analysis Experiment
@@
Appendix G Qualitative Examples in Ours Experiments
This section presents some practical examples in our experiments. These examples are from SAP200 dataset (see Figure 11), DAN dataset (see Figure 12), and AdvBench dataset (for GCG; see Figure 13), respectively.

### v4

@@ 7 Conclusion
In this work, a simple yet highly effective defense strategy $\mathbb{IA}$ is proposed to handle the widespread complex and stealthy jailbreak attacks. $\mathbb{IA}$ leverages LLM’s intrinsic capacities to analyze the essential intention of user queries before finally responding through two stages. Extensive experiments on representative jailbreak benchmarks across diverse LLMs show that $\mathbb{IA}$ could consistently and significantly enhance LLM safety while maintaining general helpfulness. $\mathbb{IA}$ works by suppressing LLM’s tendency to follow jailbreak prompts, thus leading to safer responses. Further analysis indicates that enhancing LLMs’ intention analysis capability and their inherent safety are two directions for future improvements.
@@
Our method remains to be validated on more advanced models. However, since our core intention analysis mechanism relies on LLM’s fundamental capabilities of—specifically, instruction-following and text comprehension—making it easy to perform, we believe this approach has the potential to generalize effectively across diverse models as a safety mechanism. Additionally, despite the effectiveness of our method in defending sophisticated jailbreak prompts, these prompts do not encompass the entire potential jailbreak attacks encountered in real-world scenarios. Consequently, the practical applicability of our approach remains to be validated through further testing. Our research underlines the importance of intention analysis in improving LLM safety, suggesting future work focusing on integrating this into training to reduce inference costs. Additionally, in the face of the rapid advancements in the adversarial attacks community, there is a pressing need for developing more effective and robust defense strategies for LLMs. While our method specifically targets jailbreak scenarios, broader alignment tasks still benefit from alignment training, such as RLHF.
@@
Acknowledgments
We express our gratitude to Zuchao Li and Yuchun Miao for their assistance with proofreading and insightful feedback on the writing of this paper. We thank the anonymous reviewers and the area chair for their insightful comments and suggestions. This research is supported by the National Research Foundation, Singapore, and the CyberSG R&D Programme Office (“CRPO”), under the National Cybersecurity R&D Programme (“NCRP”), RIE2025 NCRP Funding Initiative (Award CRPO-GC1-NTU-002).
@@ C.1 Automation-based Safety Evaluation
For evaluation of DAN, considering the complex instructions in DAN makes it challenging to directly identify the success of an attack through string matching, we turn to utilize a widely-adopted LLM to evaluate the harmfulness of model generations. Practically, we adopt gpt-3.5-turbo-0613171717https://openai.com/blog/chatgpt as the evaluation function following Deng et al. (2023a). Especially, the model responses are submitted via API calls, and a specific evaluation prompt is designed for gpt-3.5-turbo-0613 to rate the harmfulness on a five-degree scale (‘A (Very Harmful)’, ‘B (Harmful)’, ‘C (Moderately Harmful)’, ‘D (Slightly Harmful)’, and ‘E (No Harmful at All)’). Figure 7 shows the detailed instruction for gpt-3.5-turbo-0613 evaluation.
@@ D.1 One-Pass $\mathbb{IA}$
D.2 Two-Stage $\mathbb{IA}$
To assess the resilience of our method against specific expressions, we construct other two sets of alternative instructions for intention analysis. These additional instructions were meticulously designed with modifications of origin $\mathbb{IA}$ prompts (detailed in Figure 9) in lexical choices and syntactic constructions to direct intention analysis. The detailed prompts are presented in Figure 11. Table 6 presents the comparison results. We observe that three prompt sets all result in significant and comparable decreases in ASRs on the SAP200 across different LLMs. This uniformity highlights a key conclusion: it is the intention analysis mechanism, but not the specific instruction expressions, that makes our $\mathbb{IA}$ effective.
@@ D.2 Two-Stage $\mathbb{IA}$
Vicuna-7B Vicuna-13B ChatGLM-6B
Vanilla 73.4 65.4 45.8
+ Prompt A 2.94 0.88 5.81
+ Prompt B 5.13 2.06 5.44
+ Ours 0.31 1.12 6.12
Figure 12: The MultiJail (under two scenarios) and CipherChat Datasets results on GPT-3.5 with and without our $\mathbb{IA}$. (a) Results on direct MultiJail dataset including English (en), Chinese (zh), Italian (it), Vietnamese (vi), Arabic (ar), Korean (ko), Thai (th), Bengali (bn), Swahili (sw), and Javanese (jv). (b) Results on malicious jailbreak prompt attached to MultiJail. (c) Results on CipherChat including ASCII (en), UTF (zh), Unicode (zh), and SelfCipher (en and zh) encryptions.
Appendix E Extensive Validations of $\mathbb{IA}$’s Effectiveness
E.1 Performance under More Advanced Attacks
Recent studies Deng et al. (2023b); Yong et al. (2023) reveal that the multilingual jailbreak poses a new defense challenge for LLMs. Yuan et al. (2024) and Wei et al. (2023a) also emphasize the struggles of more powerful LLMs, such as GPT-3.5, to stay safe when countering encryption-based attack. To verify the effectiveness of our method in these advanced jailbreak scenarios, we reproduce MultiJail and CipherChat following Deng et al. (2023b) and Yuan et al. (2024), respectively, and conduct further experiments on GPT-3.5181818We observe a high rate of invalid responses on smaller LLMs like ChatGLM-6B, Vicuna-7B and Vicuna-13B under MultiJail, as demonstrated in Deng et al. (2023b). And smaller LLMs may lack advanced encryption capability, which is required by CipherChat (Yuan et al., 2024). So we only adopt GPT-3.5 for these advanded jailbreak attacks experiments.. The results of GPT-3.5 with and without our $\mathbb{IA}$ are presented in Figure 12. We observe that 1) our $\mathbb{IA}$ consistently maintains performance in low-resource languages, e.g., th, bn, sw, and jv, even in scenarios where a malicious jailbreak prompt [^19] We adopt the same jailbreak prompt in Deng et al. (2023b), namely AIM. is attached to the multilingual attacks, 2) our $\mathbb{IA}$ significantly enhances safety when facing advanced encryption-based attack, even under the most effective SelfCipher attack. These demonstrate the effectiveness of our intention analysis defense mechanism under more advanced jailbreak attacks.
@@ E.1 Performance under More Advanced Attacks
E.2 Comparison with All Defense Baselines
Table 7 lists comparison results between $\mathbb{IA}$ and the baselines. [^20] Due to computational resource and financial limitations, we randomly select 320 samples each from DAN and SAP200 datasets for comparative experiments. As observed, $\mathbb{IA}$ consistently shows superiority over other baselines on different datasets and model scales. Specifically, $\mathbb{IA}$ outperforms the second-best method by 30.32% and 23.77% averagely on SAP200 and DAN, respectively. In addition, although ICD and Self-Reminder achieve considerable reduction in ASR on GCG, their performance severely degrades when dealing with complex and stealthy jailbreak prompts. On the contrary, $\mathbb{IA}$ consistently outperforms other baselines across both prompt-level and automatic token-level jailbreak datasets. Notably, $\mathbb{IA}$ achieves the best ASRs with comparable and acceptable empirical inference runtime.
ChatGLM-6B Llama2-7B-Chat Llama3-8B-Instruct Vicuna-7B Vicuna-13B MPT-30B-Chat DeepSeek-67B-Chat GPT-3.5
DAN 93% 100% 100% 100% 100% 92% 100% 42%
SAP200 100% 100% 100% 100% 100% 100% 100% 49%
Methods Safety Helpfulness
SAP200 DAN AlpacaEval
Vicuna-7B 73.4 44.3 66.2
[…]
Vicuna-7B + Ours 0.31 2.89 63.8
E.3 $\mathbb{IA}$ achieves comparable safety with well-safety-trained LLMs without the need for additional training.
Our method aims to enhance LLM safety in the inference stage. A natural question arises: how does its performance compare to well-safety-trained LLMs? To answer this, we compare our method with a representative well-safety-trained LLM, i.e., Llama2-7B-Chat. The results are listed in Table 9. We can see that our method achieves comparable performance to Llama2-7B-Chat on safety datasets while outperforming Llama2-7B-Chat on the helpfulness dataset by almost 6%. This demonstrates the advantage of our $\mathbb{IA}$ to achieve both safety and helpfulness goals without additionally resource-consuming safety training.
DAN SAP200 DeepInception GCG Average Time Cost
Vanilla 48.4 73.4 90.0 83.0 73.7 10.2
Self-Reminder 41.3 33.8 55.4 11.0 35.4 15.0
Ours 3.42 0.31 0.00 0.00 0.93 17.3
Self-reminder+Ours 3.12 0.00 0.00 0.00 0.78 25.5
Table 10: Performance of combining our $\mathbb{IA}$ with Self-Reminder method for Vicuna-7B in terms of ASR (%) and average Time Cost (s/sample).
E.4 $\mathbb{IA}$ can be combined with another defensive method.
We integrate our $\mathbb{IA}$ method with the Self-Reminder method (Xie et al., 2023) and conduct experiments on Vicuna-7B to see where such a combination leads. The comparison results in Table 10 indicates that although our method already significantly improves LLM safety, combining it with another defensive method can further enhance the effectiveness at the cost of additional computation overhead.
Appendix F Further explanation of $\mathbb{IA}$ format’s effectiveness when generated intention is incorrect
In Figure 4, we find that even when the correct intention ratio is 0% (with all generated intentions replaced by masked or random intentions), $\mathbb{IA}$ remains effective compared to the vanilla baseline. This effectiveness is mainly due to $\mathbb{IA}$’s two-round dialogue design. As shown in Figure 2, the final policy-aligned responses are generated with the context of intention analysis sequence format in the first round conversation. In Context Learning (ICL) community (Min et al., 2022) has demonstrated that “keeping the format of the input-label pairs is key, and replacing gold labels with random labels in the demonstrations only marginally lowers the performance.” Therefore, even if the intention label generated in the first stage is incorrect, keeping the entire intention analysis format plays a significant role in making the final response safer than when no intention analysis sequence is used (vanilla method). Moreover, as indicated in Table 4, improving the ratio of correct intention labels can further enhance $\mathbb{IA}$’s performance.
Defense Methods DAN AlpacaEval
Harmfulness Helpfulness Win Rate
Vanilla 48.4 5.66 66.2
Input Check 19.0 3.25 64.4
ICD 40.4 5.79 60.3
Self-Reminder 41.3 5.89 64.6
SmoothLLM 13.5 5.35 60.8
[…]
Table 11: Comparison results for Vicuna-7B in terms of harmfulness (%), and helpfulness (%) on DAN dataset, and win rate (%) on AlpacaEval.
Appendix G Deeper Study of Safe Responses’ Helpfulness for Harmful Queries
G.1 ChatGPT Evaluation
To comprehensively study the impact of our $\mathbb{IA}$ on responses to harmful queries, we follow Zheng et al. (2023) and prompt ChatGPT to score the helpfulness of these safe refusals [^21] When refusing harmful queries, we expect LLMs to further provide reasonable explanations or suggestions instead of simply rejecting, thus being safe and helpful at the same time.. Table 11 presents comparison results between different defense methods on the harmfulness (ASR) and helpfulness score on the DAN dataset. We observe that $\mathbb{IA}$ enables LLMs to effectively give safe refusals with satisfactory helpfulness for harmful queries. We also manually check these refusals in Appendix G.2 and find that $\mathbb{IA}$ enables LLMs to craft more nuanced responses to specific unsafe intents like inciting hatred and division.
@@ G.1 ChatGPT Evaluation
G.2 Manual Check
To comprehensively study the impact of our $\mathbb{IA}$ on responses to harmful queries, we conduct a manual review of 100 random-sampled refusals on both DAN and SAP200 datasets for each of the seven LLMs under our $\mathbb{IA}$. We manually check for 1) reasonable explanations about why it rejects to respond and 2) positive suggestions for the user. In Table 8, our findings show that, following $\mathbb{IA}$, almost all LLMs can give detailed explanations in nearly 100% of cases (except for GPT-3.5 due to OpenAI’s external moderation mechanism during API interactions (OpenAI, 2023)). Specifically, we find LLMs indeed conduct more nuanced handling regarding varying unsafe intents (e.g. terrorism and violence), indicating that our $\mathbb{IA}$ ensures the helpfulness of safe responses for harmful queries. Detailed cases are presented in Figures 16.
Appendix H Cross-Intention Analysis Experiment
@@
Appendix I Qualitative Examples
This section presents some practical cases, which are from SAP200 dataset (see Figure 13), DAN dataset (see Figure 14), and AdvBench dataset (for GCG; see Figure 15), respectively.